# Logistic Multidimensional Data Analysis for Ordinal Response Variables using a Cumulative Link function

### v1

@@ Abstract
We present a multidimensional data analysis framework for the analysis of ordinal response variables. Underlying the ordinal variables, we assume a continuous latent variable, leading to cumulative logit models. The framework includes unsupervised methods, when no predictor variables are available, and supervised methods, when predictor variables are available. We distinguish between dominance variables and proximity variables, where dominance variables are analyzed using inner product models, whereas the proximity variables are analyzed using distance models. An expectation-majorization-minimization algorithm is derived for estimation of the parameters of the models. We illustrate our methodology with data from the International Social Survey Programme.
@@ 1 Introduction
In regression modeling of an ordinal response variable, the following model for the latent variable is assumed
@@
In regression modelling, the de facto default choice for the analysis of categorical response variables are logistic models (Agresti,, 2002). For binary response variables, standard binary logistic regression models have been developed and these have been extended for ordinal variables and nominal variables (see Agresti,, 2002, chapter 7). In logistic regression models for ordinal variables, we use the cumulative density function of the logistic distribution, such that $F$ equals
@@
and the corresponding regression model is known as the proportional odds model, or, more generally, the cumulative logistic regression model (McCullagh,, 1980; Anderson and Philips,, 1981; Agresti,, 2002).
@@
In classical multivariate analysis, principal component analysis (PCA, Pearson,, 1901; Hotelling,, 1936; Jolliffe,, 2002) is the standard tool for the analysis of dominance data whereas multidimensional unfolding (MDU, Heiser,, 1981; Busing,, 2010) is the standard tool for the analysis of single-peaked processes.
@@
Another multidimensional data analysis approach is multidimensional unfolding. Where PCA uses an inner-product representation of the data, MDU uses a distance representation. MDU is a generalization of multidimensional scaling to two-way two-mode data matrices (Carroll and Arabie,, 1980). In MDU, the data are approximated by the distance between so-called ideal points for the row objects (i.e., participants) and points for the response variables, that is, the distances between the two sets of points. MDU solutions can also be graphically represented by biplots, where both the row-objects and the variables are represented by points in a Euclidean space. Estimated values can be obtained by inspecting the distances between the two sets of points.
@@
PCA requires numerical data and estimation is usually done by minimizing a least squares function. Several extensions have been proposed to deal with other types of data. The first class of extensions have in common that the least squares criterion is maintained. For PCA of ordinal variables, Labovitz, (1970) proposed to use rank numbers instead of the numerical values. Korhonen and Siljamäki, (1998) proposed an ordinal PCA that maximizes the rank correlation between the original variables and the principal components. As PCA can also be conceived as an eigenvalue decomposition of the correlation matrix, Kolenikov and Angeles, (2004) argue to use an eigenvalue decomposition of the polychoric correlation matrix. Nonlinear Principal Component analysis has been proposed for the analysis of ordinal variables where an optimal transformation of the original variables is sought that minimizes the least squares loss function (Gifi,, 1990; Linting et al.,, 2007; Mori et al.,, 2016; Hoshiyar et al.,, 2021). The second class of extensions use other loss functions. For binary data, several authors (Schein et al.,, 2003; De Leeuw,, 2006; Landgraf and Lee,, 2020) proposed PCA using the binomial negative log-likelihood as loss function. Collins et al., (2001) proposed a generalization of PCA to the exponential family to deal with, for example, binary data or integer-valued data such as count data. As far as we know, there are no exponential family generalizations of PCA for ordinal response variables.
@@
For constrained PCA, that is reduced rank regression or redundancy analysis, generalizations have been proposed for binary data (De Rooij,, 2023) and the exponential family (Yee and Hastie,, 2003). As far as we know, there are no exponential family generalizations of these reduced rank models for ordinal response variables.
@@
MDU usually requires two-way two-mode data with non-negative values that can be interpreted as proximities. Metric MDU requires numerical proximity data. Non-metric MDU can deal with, for example, rank order data, but has been plagued by degeneracy problems for ages (for an overview see Busing, (2010), chapter 1). MDU has been generalized for binary variables by several authors (Andrich,, 1988; Takane,, 1998; DeSarbo and Hoffman,, 1986), using squared distances. Andrich, (1988) proposed a unidimensional model that does not allow for predictor variables. Takane, (1998) and DeSarbo and Hoffman, (1986) describe generalizations to multiple dimensions and explanatory versions. De Rooij et al., (2022) defined a model on the basis of (unsquared) distances for binary data, both with and without predictor variables. As far as we know, there are no exponential family generalizations of MDU for ordinal response variables.
@@
In this manuscript, we will propose a family of geometric, multidimensional, models for multivariate ordinal data. We distinguish between dominance and proximity response variables. We also distinguish models with and without predictor variables. One unified algorithm for the estimation of model parameters will be presented (Section 3), where in the lower level iterations updates differ between the four approaches. We will also briefly discuss model selection issues. In Section 4, we show applications to data from the International Social Survey Programme (ISSP Research Group,, 2022). We end this paper with some conclusions.
@@ 2 Cumulative Logistic Multidimensional Models
We consider a set of ordinal variables $y_{ir}$ ($i=1,\ldots,N$, $r=1,\ldots,R$) where variable $r$ has $C_{r}$ categories, coded as $c=1,\ldots,C_{r}$. Underlying each ordered categorical response variable $y_{r}$ we assume a continuous latent variable $z_{r}$. We model these latent variables as
@@
When predictor variables are available for the participants, the coordinates of the principal scores or ideal points can be restricted to be a linear, additive function of these predictor variables, that is $\bm{u}_{i}=\bm{B}^{\prime}\bm{x}_{i}$ with $\bm{B}$ a $P\times S$ matrix. Within the PCA context, we obtain the reduced rank regression model. For MDU, we obtain a restricted multidimensional unfolding.
@@
where, similar to the proportional odds regression model, the thresholds are category specific but the structural part of the model is variable specific.
@@ 2.1 Properties of Cumulative Logistic Models
Let us consider two subjects with locations $\bm{u}_{1}$ and $\bm{u}_{2}$. The cumulative log-odds ratio for response variable $r$ is
@@
With a PCA parameterisation, $\tau$ can be written as
@@
which does not depend on $c$. This is the proportional odds assumption. Furthermore, if we define $\bm{u}_{2}$ as $\bm{u}_{2}=\bm{u}_{1}+\delta(\bm{u})$, such that $\delta(\bm{u})$ is a shift from one position to another, then we may write
@@
which shows that it does not matter where in the Euclidean space this shift happens, the cumulative log-odds ratio remains constant for constant $\delta(\bm{u})$.
When $\bm{u}_{i}=\bm{B}^{\prime}\bm{x}_{i}$, we might be interested in the change of cumulative log-odds when one of the predictor variables increases by one unit. Say, the $p$-th predictor variable increases by a unit, such that, $\delta(\bm{u})=\bm{B}^{\prime}(\bm{x}_{2}-\bm{x}_{1})=\bm{b}_{p}$, then the cumulative log-odds increase by $\left\langle\bm{b}_{p},\bm{v}_{r}\right\rangle$.
Cumulative logistic reduced rank regression models can also be interpreted numerically, similar to the regression weights in proportional odds models. With reduced rank coefficient matrix $\bm{BV}^{\prime}$, each column of this matrix represents a change in cumulative log odds for a predictors unit increase.
For cumulative logistic MDU there is a nonlinear, distance relationship. Therefore, in contrast with the PCA parameterization, there is no such thing as a proportional odds assumption. The changes in cumulative log odds are not constant for changes in $\bm{u}$.
@@ 2.2 Model Selection
For the AIC and BIC we need the number of parameters. In all our models, we have the threshold parameters, the number of which is $\sum_{r}(C_{r}-1)$. For reduced rank models, such as PCA and RRR, the number of parameters in the structural part is $(N+R-S)S$ and $(P+R-S)S$, respectively. Mukherjee et al., (2015) showed, in the context of linear reduced rank models, that these numbers are naive estimates, For linear reduced rank model better estimates are available, but this theory is not extended (yet) to other types of models, such as ordinal response variables. For MDU and RMDU, the number of parameters in the structural part is $(N+R)S-S(S+1)/2$ and $(P+R)S-S(S-1)/2$, respectively.
@@ 2.3.1 CLPCA biplots
Biplots (Gower and Hand,, 1996; Gower et al.,, 2011) are useful displays for the results of a PCA, especially for two-dimensional solutions. We will now discuss the geometry of the two-dimensional biplot for Cumulative Logistic PCA. Like a usual PCA biplot, observations are shown as points, and variables are shown by axes. The coordinates of the points are given by the estimated $\bm{u}_{i}$. The variable axes are straight lines though the origin with direction $\frac{v_{r2}}{v_{r1}}$. In Figure 2a, we present a simplified biplot where the observations are shown by grey dots and there is a single variable axis (solid line). For this variable, $\bm{v}_{r}=[1,0.5]^{\prime}$ and suppose $m_{r_{c}}=(-2.0,-1.5,-0.5)$ for $c=1,2$ and $3$, respectively, i.e., thresholds for a four-point response scale. Estimated values for the response of an individual can be obtained by projecting the point representing this individual onto the variable axis. Subjects positioned in the lower left corner have lower expected values for the response, while subjects in the upper right corner have higher expected values, projecting higher onto the variable axis. To further increase interpretation and provide numerical values for the expected value, Gower and Hand, (1996) suggest to add labeled markers to the variable axis. For CLPCA there are several possibilities.
@@
This inner product is constant ($\mu$, say) for all points on a line projecting at the same location of the variable axis. Therefore, the point of projection may be calibrated by labelling this point with the value $\mu$. This value also applies to the point of projection itself, which is $\lambda\bm{v}_{r}$. For the point $\lambda\bm{v}_{r}$ to be calibrated $\mu$, it must satisfy
@@
so that $\lambda=\mu/\bm{v}_{r}^{\prime}\bm{v}_{r}$ and the coordinates of the point on the variable axis that is calibrated with a value of $\mu$ are $\mu\bm{v}_{r}/(\bm{v}_{r}^{\prime}\bm{v}_{r})$. As such, we would have the markers expressing values of the underlying latent continuous variable, but the interest lies in the observed ordinal response variable. The estimated response is $\hat{y}=c$ if $m_{r_{c-1}}\leq\hat{z}<m_{r_{c}}$. Therefore, markers can be based on the estimated thresholds (Anderson and Philips,, 1981). These markers indicate the transition points between adjacent categories. The coordinates of the marker point are given by $\hat{m}_{r_{c}}\bm{v}_{r}/(\bm{v}_{r}^{\prime}\bm{v}_{r})$ and these can be labeled by 1|2, 2|3, and so forth. The application of these markers is illustrated in Figure 2a. An advantage of these markers over the ones based on posterior probabilities is that each threshold is represented. The two-dimensional space can be partitioned into $C_{r}$ areas by drawing $C_{r}-1$ decision lines orthogonal on the variable axis and through the marker points, these are represented by the dotted lines in Figure 2a. In an usual biplot, several response variables are represented by variable axes jointly partitioning the two-dimensional space in open and closed regions that each represent a particular response profile.

### v2

@@ Abstract
We present a multidimensional data analysis framework for the analysis of ordinal response variables. Underlying the ordinal variables, we assume a continuous latent variable, leading to cumulative logit models. The framework includes unsupervised methods, when no predictor variables are available, and supervised methods, when predictor variables are available. We distinguish between dominance variables and proximity variables, where dominance variables are analyzed using inner product models, whereas the proximity variables are analyzed using distance models. An expectation-majorization-minimization algorithm is derived for estimation of the parameters of the models. We illustrate our methodology with three empirical data sets highlighting the advantages of the proposed framework. A simulation study is conducted to evaluate the performance of the algorithm.
@@ 1 Introduction
In regression modeling of an ordinal response variable, the following model for the latent variable is assumed (Anderson and Philips,, 1981)
@@
In regression modelling, the de facto default choice for the analysis of categorical response variables are logistic models (Agresti,, 2002). For binary response variables, standard binary logistic regression models have been developed and these have been extended for ordinal variables and nominal variables (see Agresti,, 2002, chapter 7). Logistic models have the advantage that detailed interpretation in terms of changes in log-odds is possible. Such an interpretation is not available for, for example, probit models that use the cumulative density of the normal distribution. Otherwise, the fit of logit and probit models is usually very similar (Agresti,, 2002). In logistic regression models for ordinal variables, we use the cumulative density function of the logistic distribution, such that $F$ equals
@@
and the corresponding regression model is known as the proportional odds model, or, more generally, the cumulative logistic regression model (Walker and Duncan,, 1967; McCullagh,, 1980; Anderson and Philips,, 1981; Agresti,, 2002).
@@
In classical multivariate analysis, principal component analysis (PCA, Pearson,, 1901; Hotelling,, 1936; Jolliffe,, 2002) is the standard multidimensional data analysis tool for the analysis of dominance variables whereas multidimensional unfolding (MDU, Heiser,, 1981; Busing,, 2010) is the standard tool for the analysis of proximity variables.
@@
Another multidimensional data analysis approach is multidimensional unfolding. MDU is targeted towards proximity variables. Where PCA uses an inner-product representation of the data, MDU uses a distance representation. MDU is a generalization of multidimensional scaling to two-way two-mode data matrices (Carroll and Arabie,, 1980). In MDU, the data are approximated by the distance between so-called ideal points for the row objects (i.e., participants) and points for the response variables, that is, the distances between the two sets of points. MDU solutions can also be graphically represented by biplots, where both the row-objects and the variables are represented by points in a Euclidean space. Estimated values can be obtained by inspecting the distances between the two sets of points.
@@
PCA, MDU and its constrained versions, RRR and RMDU, are usually estimated by least squares methods. For categorical response variables, however, linear models estimated with least squares are not optimal and might lead to distorted effect sizes, type-I errors and inversion of effects (Liddell and Kruschke,, 2018). Logistic models estimated using maximum likelihood offer an alternative.
For binary data, several authors (Schein et al.,, 2003; De Leeuw,, 2006; Landgraf and Lee,, 2020) proposed PCA using the binomial negative log-likelihood as loss function. Collins et al., (2001) proposed a generalization of PCA to the exponential family to deal with, for example, binary data or integer-valued data such as count data. As far as we know, only Vicente-Villardón and Sánchez, (2014) investigated an exponential family generalizations of PCA for ordinal response variables including a biplot visualization.
@@
For constrained PCA, that is reduced rank regression or redundancy analysis, a logistic model has been proposed for binary data by De Rooij, (2023) and (Yee and Hastie,, 2003) proposed a wider class of models for the exponential family, similar to generalized linear models (McCullagh and Nelder,, 1989). As far as we know, there are no exponential family generalizations of these reduced rank models for ordinal response variables.
@@
For MDU, several attempts can be found in the literature to exponential family multidimensional unfolding models. Andrich, (1988); Takane, (1998); DeSarbo and Hoffman, (1986) defined MDU models for binary variables using squared distances. Andrich, (1988) proposed a unidimensional model that does not allow for predictor variables. Takane, (1998) and DeSarbo and Hoffman, (1986) describe generalizations to multiple dimensions that can include predictors. De Rooij et al., (2024) defined a model on the basis of (unsquared) distances for binary data, both with and without predictor variables. As far as we know, there are no exponential family generalizations of MDU for ordinal response variables.
In conclusion, there have been several attempts to define exponential family models for PCA and MDU and their constraint versions that include predictor variables. These attempts mainly focus on binary response variables, but some include also other types of variables (Collins et al.,, 2001; Yee and Hastie,, 2003) like count variables. However, no exponential family generalizations exist for PCA or MDU of ordinal responses, neither with or without predictor variables.
The goal of this paper is to fill this gap and propose multidimensional models for ordinal response variables in the exponential family. We will develop models for multivariate ordinal dominance variables (i.e., PCA and RRR) and proximity variables (i.e., MDU and RMDU). Both models without predictor variables (i.e., PCA and MDU) and with predictor variables (i.e., RRR and RMDU) will be presented. Along the algebraic formulation, we will also develop biplot methodology for visualization of the models. One unified algorithm for maximum likelihood estimation of model parameters will be developed and tested, where in the lower level iterations updates differ between the four approaches. To illustrate the multidimensional models and the difference between the dominance and proximity perspective, we will apply the models on several data sets. In the first application, we have cognitive data and use the dominance approach. The second application highlights the difference between the dominance and proximity approaches using behavioural response variables. The third application shows in detail the proximity approach where the response variables concern attitudes concerning the environment. Biplots for all three applications will be discussed in detail.
@@
The outline of this manuscript is as follows. In Section 2, we will propose a family of geometric, multidimensional, models for multivariate ordinal data. We distinguish between dominance and proximity response variables and between models with and without predictor variables. Properties of the model will be derived. We briefly discuss model selection and discuss in detail the visualization of the models using biplots. In Section 3, a unified algorithm for maximum likelihood estimation of model parameters will be presented. In Section 4, we show three applications. The first application considers data from students performance on a statistics exam. The second and third data set are from the International Social Survey Programme (ISSP Research Group,, 2022). We test our algorithm using simulated data in Section 5. We end this paper with some discussion and conclusions.
@@ 2 Cumulative Logistic Multidimensional Models
We consider a set of ordinal variables with observed values $y_{ir}$ ($i=1,\ldots,N$, $r=1,\ldots,R$) where variable $r$ has $C_{r}$ categories, coded as $c=1,\ldots,C_{r}$. Underlying each ordered categorical response variable $y_{r}$ we assume a continuous latent variable $z_{r}$. We model these latent variables as
@@
When predictor variables are available for the participants, the coordinates of the principal scores or ideal points can be restricted to be a linear, additive function of these predictor variables, that is $\bm{u}_{i}=\bm{B}^{\prime}\bm{x}_{i}$ with $\bm{B}$ a $P\times S$ matrix. Within the PCA context, we obtain the reduced rank regression model (CLRRR). For MDU, we obtain a restricted multidimensional unfolding (CLRMDU).
@@
where, similar to the proportional odds regression model, the thresholds ($m_{r_{c}}$) are category specific but the structural part ($\theta_{ir}$) of the model is variable specific.
In this section, we introduced a framework of four models for multidimensional data analysis of ordinal response variables. We first distinguished between models for dominance and proximity variables. Within each of these two types, we distinguish between models with or without predictor variables.
@@ 2.1 Properties of Cumulative Logistic Models
Let us consider two subjects with locations $\bm{u}_{1}$ and $\bm{u}_{2}$. The cumulative log-odds ratio for response variable $r$ is defined as
@@
With a PCA (or RRR) parameterisation, $\tau$ can be written as
@@
which does not depend on $c$. This shows that CLPCA (and CLRRR) make a proportional odds assumption, for a given change in the positions for the subjects all the cumulative log-odds for variable $r$ change with the same amount. Furthermore, if we define $\bm{u}_{2}$ as $\bm{u}_{2}=\bm{u}_{1}+\delta(\bm{u})$, such that $\delta(\bm{u})$ is a shift from one position to another, then we may write
@@
which shows that it does not matter where in the Euclidean space this shift happens, the cumulative log-odds ratio remains constant for constant $\delta(\bm{u})$. Consider participants 1 and 2, with coordinates $\bm{u}_{1}=(0,0)$ and $\bm{u}_{2}=(1,0.5)$. The estimated cumulative log-odds ratio for these two participants is the same as for participants 3 and 4 with coordinates $\bm{u}_{3}=(-5,3)$ and $\bm{u}_{4}=(-4,3.5)$, as the difference between these pairs of coordinates is the same.
When predictor variables are used in the analysis, we constrain the coordinates to be linear combinations of those variables, that is $\bm{u}_{i}=\bm{B}^{\prime}\bm{x}_{i}$. Of interest is then the comparison between two persons that differ one unit in one of the predictor variables and have equal values for the other predictors. Say, the $p$-th predictor variable increases by a unit, such that, $\delta(\bm{u})=\bm{B}^{\prime}(\bm{x}_{2}-\bm{x}_{1})=\bm{b}_{p}$, then the cumulative log-odds increase by $\left\langle\bm{b}_{p},\bm{v}_{r}\right\rangle$.
Below in Section 2.3, we will describe biplots for the interpretation of our multidimensional models. Cumulative logistic reduced rank regression models can also be interpreted numerically, similar to the regression weights in a proportional odds model. With reduced rank coefficient matrix $\bm{A}=\bm{BV}^{\prime}$, each column of this matrix represents a change in cumulative log odds for the corresponding response for unit increase in the predictor.
For cumulative logistic MDU and restricted MDU, there is a nonlinear, distance relationship. The cumulative log-odds ratio becomes
$\tau=d(\bm{u}_{1},\bm{v}_{r})-d(\bm{u}_{2},\bm{v}_{r}),$
again not depending on $c$, only on the distances. This is again the proportional odds assumption, but in contrast with the PCA parameterization, however, the changes in cumulative log odds are not constant for changes in $\bm{u}$. That is, using the example with participants 1, 2, 3, and 4, described above again, the cumulative log-odds ratio for participants 1 and 2 is not equal to that of participants 3 and 4. Similarly, unit changes in one of the predictors do not lead to a constant change in cumulative log-odds (i.e., the estimated $\tau$ is different for two participants with predictor values 0 and 1 compared to two participants with predictor values 2 and 3, say). Furthermore, unit changes in one of the predictors (say, $x_{1}$) lead to different changes in $\tau$ for participants with varying values on the other predictors. Although the relationship of the predictors is additive in defining the positions of the participants in the biplot, the relationship between predictors is not additive when looking at the effect on the response variables. Results can therefore not be represented numerically and we have to rely on the biplot visualizations described in detail below.
@@ 2.2 Model Selection
For the AIC and BIC we need the number of parameters. In all our models, we have the threshold parameters, the number of which is $\sum_{r}(C_{r}-1)$. For PCA and RRR, the number of parameters in the structural part is $(N+R-S)S$ and $(P+R-S)S$, respectively. Mukherjee et al., (2015) showed, in the context of linear reduced rank models, that these numbers are naive estimates. For linear reduced rank model better estimates are available, but this theory is not extended (yet) to other types of models, such as ordinal response variables. For MDU and RMDU, the number of parameters in the structural part is $(N+R)S-S(S-1)/2$ and $(P+R)S-S(S-1)/2$, respectively.
@@ 2.3 Biplots
In this section, we will discuss biplots for the visualization of the models results. These biplots are most valuable for two-dimensional solutions, but can also be used for visualizaztion of pairs of dimensions in case of higher dimensional solutions. We first discuss biplots for CLPCA and CLMDU. Afterwards, we discuss the case when predictor variables are available for the analysis. In that case, the biplots are extended with extra information about the predictor variables.
@@ 2.3.1 CLPCA biplots
Biplots (Gabriel,, 1971; Gower and Hand,, 1996; Gower et al.,, 2011) are useful displays for the results of a PCA, especially for two-dimensional solutions. We will now discuss the geometry of the two-dimensional biplot for Cumulative Logistic PCA. Like a usual PCA biplot, observations are shown as points, and variables are shown by axes. The coordinates of the points are given by the estimated $\bm{u}_{i}$. The variable axes are straight lines though the origin with direction $\frac{v_{r2}}{v_{r1}}$. In Figure 2a, we present a simplified biplot where the observations are shown by grey dots and there is a single variable axis (solid line). For this variable, $\bm{v}_{r}=[1,0.5]^{\prime}$ and suppose $m_{r_{c}}=(-2.0,-1.5,-0.5)$ for $c=1,2$ and $3$, respectively, i.e., thresholds for a four-point response scale. Estimated values for the response of an individual can be obtained by projecting the point representing this individual onto the variable axis. Subjects positioned in the lower left corner have lower expected values for the response, while subjects in the upper right corner have higher expected values, projecting higher onto the variable axis. To further increase interpretation and provide numerical values for the expected value, Gower and Hand, (1996) suggest to add labeled markers to the variable axis. For CLPCA there are several possibilities.
@@
This inner product is constant ($\mu$, say) for all points on a line projecting at the same location of the variable axis, that is, a line orthogonal to the variable axis. Therefore, the point of projection may be calibrated by labelling this point with the value $\mu$. This value also applies to the point of projection itself, which is $\lambda\bm{v}_{r}$. For the point $\lambda\bm{v}_{r}$ to be calibrated $\mu$, it must satisfy
@@
so that $\lambda=\mu/\bm{v}_{r}^{\prime}\bm{v}_{r}$ and the coordinates of the point on the variable axis that is calibrated with a value of $\mu$ are $\mu\bm{v}_{r}/(\bm{v}_{r}^{\prime}\bm{v}_{r})$. As such, we would have the markers expressing values of the underlying latent continuous variable, but the interest lies in the observed ordinal response variable. The estimated response is $\hat{y}=c$ if $m_{r_{c-1}}\leq\hat{z}<m_{r_{c}}$. Therefore, markers can be based on the estimated thresholds. These markers indicate the transition points between adjacent categories. The coordinates of the marker point are given by $\hat{m}_{r_{c}}\bm{v}_{r}/(\bm{v}_{r}^{\prime}\bm{v}_{r})$ and these can be labeled by 1|2, 2|3, and so forth. The application of these markers is illustrated in Figure 2a. An advantage of these markers over the ones based on posterior probabilities is that each threshold is represented.
Based on these markers, the two-dimensional space can be partitioned into $C_{r}$ areas by drawing $C_{r}-1$ decision lines orthogonal on the variable axis and through the marker points, these are represented by the dotted lines in Figure 2a. In an usual biplot, several response variables are represented by variable axes jointly partitioning the two-dimensional space in open and closed regions that each represent a particular response profile.